### v1

# References

[…]
Acknowledgements
This work was done while visiting the Academia Sinica, Taipei for the Pacific-Rim Graph Theory Group Workshop in 2024 organized by Bruce Reed. Tony Huynh thanks Stefan Weltge from whom he learned about the tea sharing problem at the 26th Combinatorial Optimization Workshop in Aussois, France in 2024. We are also grateful to Nina Gantert and Timo Vilkas for providing valuable references and background information.
J. Pascal Gollin was supported by the Insitute for Basic Science (IBS-R029-Y3). Kevin Hendrey and Sang-il Oum were supported by the Insitute for Basic Science (IBS-R029-C1). Hao Huang was supported in part by a start-up grant at NUS and an MOE Academic Research Fund (AcRF) Tier 1 grant of Singapore. Tony Huynh was partially supported by Progetti Grandi di Ateneo Programme, Grant RG123188B3F7414A (ASGARD). Bojan Mohar was supported through the ERC Synergy grant KARST (European Union, ERC, KARST, project number 101071836). Wei-Hsuan Yu was supported by MOST of Taiwan under Grant109-2628-M-008-002-MY4. Xuding Zhu was supported by National Science Foundation of China with grant numbers NSFC 12371359 and U20A2068.
[…]
Appendix A Proposition 2.1 with general sharing moves
We present a direct proof of Corollary 1.2 (without needing to take a limit). This proof is adapted from our proof of Proposition 2.1.
An integer interval is a set of the form ${[a,b]}$, for some integers $a$ and $b$. Let $G$ be a graph and let ${r\in V(G)}$. Note that if $T$ is a finite nonempty subset of ${V(G)}$ which induces a connected subgraph, then the set ${\{d_{G}(r,v)\mid v\in T\}}$ is an integer interval. We start our first proof by an inequality on integer intervals.
Lemma A.1.
Let $s$, $t$ be non-negative integers such that $s+t>0$. Let $x_{1}$, $x_{2}$, $\ldots$, $x_{s}$, $y_{1}$, $y_{2}$, $\ldots$, $y_{t}$ be non-negative integers such that ${\{x_{1},x_{2},\ldots,x_{s},y_{1},y_{2},\ldots,y_{t}\}}$ is an integer interval. Then
$\frac{s}{s+t}\prod_{i=1}^{s}x_{i}\prod_{j=1}^{t}y_{i}+\frac{t}{s+t}\prod_{i=1}^{s}(x_{i}+1)\prod_{j=1}^{t}(y_{j}+1)\geq\prod_{i=1}^{s}x_{i}\prod_{j=1}^{t}(y_{j}+1).$
Proof.
We may assume that $s,t>0$. Since permuting values of ${(x_{1},\ldots,x_{s},y_{1},\ldots,y_{t})}$ does not change the left-hand side, we may assume that ${y_{t}\leq\cdots\leq y_{1}\leq x_{1}\leq\cdots\leq x_{s}}$ to maximize the right-hand side. We may also assume that ${x_{i}>0}$ for all $i$. By dividing both sides by ${\prod_{i=1}^{s}x_{i}\prod_{j=1}^{t}(y_{j}+1)}$, it is enough to prove that
$\frac{s}{s+t}\prod_{j=1}^{t}\frac{y_{i}}{y_{j}+1}+\frac{t}{s+t}\prod_{i=1}^{s}\frac{x_{i}+1}{x_{i}}\geq 1.$
Since ${\frac{1+x}{x}}$ is a decreasing function for ${x>0}$, we may assume that ${x_{i}=x_{1}+i-1}$ by increasing values of $x_{i}$ whenever possible, as long as ${\{x_{1},x_{2},\ldots,x_{s},y_{1},y_{2},\ldots,y_{t}\}}$ is an integer interval. Therefore,
$\prod_{i=1}^{s}\frac{x_{i}+1}{x_{i}}=\frac{x_{1}+s}{x_{1}}.$
If ${x_{1}\leq t}$, then ${\frac{t}{s+t}\frac{x_{1}+s}{x_{1}}\geq 1}$. Therefore we may assume that ${x_{1}>t}$.
Since ${\frac{y}{1+y}}$ is an increasing function for ${y>0}$, we may assume that ${y_{i}=x_{1}-i}$ by decreasing values of $y_{i}$ whenever possible, as long as ${\{x_{1},x_{2},\ldots,x_{s},y_{1},y_{2},\ldots,y_{t}\}}$ is an integer interval. Then
$\prod_{i=1}^{t}\frac{y_{i}}{y_{i}+1}=\frac{x_{1}-t}{x_{1}}.$
Now we deduce that $\frac{s}{s+t}\frac{x_{1}-t}{x_{1}}+\frac{t}{s+t}\frac{x_{1}+s}{x_{1}}=1$. This proves the inequality. ∎
We are now ready to prove the following strengthened version of Proposition 2.1.
Proposition A.2.
Let $G$ be a connected graph, let ${r\in V(G)}$, and let ${w_{0}\in\mathbb{R}_{\geq 0}^{V(G)}}$ be $r$-feasible. Then every ${w\in\mathbb{R}_{\geq 0}^{V(G)}}$ on $G$ that is reachable from $w_{0}$ is $r$-feasible.
Proof.
By normalizing, we may assume that ${w_{0}(G)=1}$. Let ${(w_{0},\dots,w_{t})}$ be a sequence of weight distributions on $G$ such that ${w_{t}=w}$ and $w_{i}$ is obtained from $w_{i-1}$ by a single sharing move for each ${i\in[t]}$. We proceed by induction on $t$. We may assume that ${t\geq 1}$; otherwise there is nothing to prove. Let ${w^{\prime}=w_{t-1}}$. By induction, $w^{\prime}$ is $r$-feasible. Suppose that $w$ is obtained from $w^{\prime}$ by sharing on a finite nonempty set ${T\subseteq V(G)}$.
We may assume that ${r\notin S}$. Let ${s={\left\lvert{S\cap T}\right\rvert}}$ and ${t={\left\lvert{T-S}\right\rvert}}$. By the definition of a sharing move, ${w(S)=\frac{s}{s+t}w^{\prime}(T)+w^{\prime}(S-T)=\frac{s}{s+t}w^{\prime}(S\cup T)+\frac{t}{s+t}w^{\prime}(S-T)}$. For a finite set ${X\subseteq V(G)}$, recall ${\rho(X)=\prod_{x\in X}\frac{d_{G}(r,x)}{d_{G}(r,x)+1}}$. Since $w^{\prime}$ satisfies (2.1),
$0\leq w^{\prime}(S\cup T)\leq 1-\rho(S\cup T)\qquad\text{and}\qquad 0\leq w^{\prime}(S-T)\leq 1-\rho(S-T).$
We deduce that
$w(S)$ $\leq\frac{s}{s+t}\left(1-\rho(S\cup T)\right)+\frac{t}{s+t}\left(1-\rho(S-T)\right)$
$=1-\left(\frac{s}{s+t}\rho(S\cup T)+\frac{t}{s+t}\rho(S-T)\right)$
$=1-\rho(S)\left(\frac{s}{s+t}\rho(T-S)+\frac{t}{s+t}\frac{1}{\rho(T\cap S)}\right)$
$\leq 1-\rho(S),$
where the last inequality follows from Lemma A.1. Thus, $w$ is also $r$-feasible. ∎
Appendix B Lemma 3.4 with general sharing moves
We now present a stronger variant of Lemma 3.4 with general sharing moves. This will allow us to prove Proposition 3.5 without needing to take a limit. Throughout this appendix, $G$ is a fixed finite connected graph with $n$ vertices and $v$ is a fixed vertex of $G$. We use the same notation as in Section 3.
Lemma B.1.
Let $t\geq 2$ be an integer and let $d_{1},d_{2},\ldots,d_{t}$ be positive integers such that $\{d_{1},d_{2},\ldots,d_{t}\}$ is an integer interval. Let $w_{1},w_{2},\ldots,w_{t}$ be non-negative reals such that $w_{i}\leq w_{i+1}$ for all $i\in[t-1]$. Then
$\sum_{i=1}^{t}\left(\prod_{j=1}^{i-1}(1+d_{j})\right)\left(\prod_{k=i+1}^{t}d_{k}\right)w_{i}\geq\left(\sum_{i=1}^{t}\left(\prod_{j=1}^{i-1}(1+d_{j})\right)\left(\prod_{k=i+1}^{t}d_{k}\right)\right)\frac{1}{t}\sum_{i=1}^{t}w_{i}.$
Proof.
First, it is easy to see that $\sum_{i=1}^{t}\prod_{j=1}^{i-1}(1+d_{j})\prod_{k=i+1}^{t}d_{k}=\prod_{j=1}^{t}(1+d_{j})-\prod_{k=1}^{t}d_{k}$ and therefore permuting the values of $(d_{1},d_{2},\ldots,d_{t})$ does not change the right-hand side of the inequality.
We claim that we may assume that $d_{\ell}\leq d_{\ell+1}$ for all $\ell\in[t-1]$. Suppose $d_{\ell}>d_{\ell+1}$ for some $\ell\in[t-1]$. Observe that
$(d_{\ell+1}w_{\ell}+(1+d_{\ell})w_{\ell+1})-(d_{\ell}w_{\ell}+(1+d_{\ell+1})w_{\ell+1})=(d_{\ell+1}-d_{\ell})(w_{\ell}-w_{\ell+1})\geq 0$
and therefore swapping the values of $d_{\ell}$ and $d_{\ell+1}$ does not increase the left-hand side. This proves the claim.
Now observe that $\left(\prod_{j=1}^{i-1}(1+d_{j})\prod_{k=i+1}^{t}d_{k}\right)\leq\left(\prod_{j=1}^{i}(1+d_{j})\prod_{k=i+2}^{t}d_{k}\right)$ for each $i\in[t-1]$, because $d_{i+1}\leq 1+d_{i}$. By Chebyshev’s sum inequality, we deduce the desired inequality. ∎
We are now ready to prove the following strengthened version of Lemma 3.4.
Lemma B.2.
Let ${w\in\mathbb{R}_{\geq 0}^{V(G)}}$ and let $w^{\prime}$ be the weight distribution on $G$ obtained from $w$ by sharing on a finite connected set $T$ of vertices. Then ${\phi_{w^{\prime}}(w^{\prime}(v))\leq\phi_{w}(w(v))}$.
Proof.
Let ${\sigma=(v_{1},\ldots,v_{n-1})}$ be an enumeration of $V(G)\smallsetminus\{v\}$ such that ${w^{\prime}(v_{i})\leq w^{\prime}(v_{i+1})}$ for all ${i\in[n-2]}$. By Lemma 3.3, ${\phi_{w^{\prime}}(w^{\prime}(v))=\phi_{\ell,w^{\prime}}(\sigma,w^{\prime}(v))}$, where
$\ell\coloneq 1+{\left\lvert{\{i\in[n-1]\mid w^{\prime}(v_{i})<w^{\prime}(v)\}}\right\rvert}.$
We may assume that ${\left\lvert{T}\right\rvert}>1$.
First, consider the case ${v\notin T}$. Then ${w(v)=w^{\prime}(v)}$. Since ${w^{\prime}(x)=\frac{w(T)}{{\left\lvert{T}\right\rvert}}}$ for every vertex $x\in T$, we may assume that ${T=\{v_{s},v_{s+1},\ldots,v_{s+t-1}\}}$ for some ${s\in[n-t]}$ with $t={\left\lvert{T}\right\rvert}$. We may also assume that ${w(v_{s+j-1})\leq w(v_{s+j})}$ for all $j\in[t-1]$. We will show that ${\phi_{\ell,w^{\prime}}(\sigma,w(v))\leq\phi_{\ell,w}(\sigma,w(v))}$. If ${s<\ell}$, then since ${w(v_{j})=w^{\prime}(v_{j})}$ for all ${j\geq\ell}$, it follows that ${\phi_{\ell,w^{\prime}}(\sigma,w(v))=\phi_{\ell,w}(\sigma,w(v))}$. So we may assume that ${\ell\leq s}$. Let
$a\coloneqq f_{(v_{\ell},v_{\ell+1},\ldots,v_{s-1});w}(w(v))=f_{(v_{\ell},v_{\ell+1},\ldots,v_{s-1});w^{\prime}}(w^{\prime}(v)).$
It suffices to show that ${f_{(v_{s},v_{s+1},\ldots,v_{s+t-1});w^{\prime}}(a)\leq f_{(v_{s},v_{s+1},\ldots,v_{s+t-1});w}(a)}$. For each $j\in[t]$, let $d_{j}\coloneqq d_{G}(v,v_{s+j-1})$ and $w_{j}\coloneqq w(v_{s+j-1})$. Then, by Lemma B.1, we have
$f_{(v_{s},v_{s+1},\ldots,v_{s+t-1});w}(a)$ $=\left(\prod\limits_{i=1}^{t}\frac{d_{i}}{d_{i}+1}\right)a+\sum_{i=1}^{t}\left(\prod\limits_{j=i+1}^{t}\frac{d_{j}}{d_{j}+1}\right)\frac{1}{d_{i}+1}w_{i}$
$=\left(\prod\limits_{i=1}^{t}\frac{d_{i}}{d_{i}+1}\right)a+\frac{\sum\limits_{i=1}^{t}\left(\prod\limits_{j=1}^{i-1}(1+d_{j})\right)\left(\prod\limits_{k=i+1}^{t}d_{k}\right)w_{i}}{\prod\limits_{i=1}^{t}(d_{i}+1)}$
$\geq\left(\prod\limits_{i=1}^{t}\frac{d_{i}}{d_{i}+1}\right)a+\frac{\sum\limits_{i=1}^{t}\left(\prod\limits_{j=1}^{i-1}(1+d_{j})\right)\left(\prod\limits_{k=i+1}^{t}d_{k}\right)}{\prod\limits_{i=1}^{t}(d_{i}+1)}\frac{\sum\limits_{i=1}^{t}w_{i}}{t}$
$=f_{(v_{s},v_{s+1},\ldots,v_{s+t-1});w^{\prime}}(a).$
So the only case left to consider is when ${v\in T}$. Since ${w^{\prime}(x)=\frac{w(T)}{{\left\lvert{T}\right\rvert}}}$ for every vertex $x\in T$, we may assume that ${T=\{v,v_{s},v_{s+1},\ldots,v_{s+t-2}\}}$ for some ${s\in[n-t+1]}$ with $t={\left\lvert{T}\right\rvert}$. We may also assume that ${w(v_{s+j-1})\leq w(v_{s+j})}$ for all $j\in[t-2]$. By Lemma 3.3,
$\phi_{w^{\prime}}(w^{\prime}(v))$ $=\phi_{s+t-1,w^{\prime}}(\sigma,w^{\prime}(v))$
$=f_{(v_{s+t-1},v_{s+t-2},\ldots,v_{n-1});w^{\prime}}(w^{\prime}(v))$
$=f_{(v_{s+t-1},v_{s+t-2},\ldots,v_{n-1});w}(w^{\prime}(v))=\phi_{s+t-1,w}(\sigma,w^{\prime}(v)).$
Since $\phi_{s+t-1,w}$ is strictly increasing, we may assume that $w(v)<w^{\prime}(v)$, because otherwise $\phi_{w^{\prime}}(w^{\prime}(v))=\phi_{s+t-1,w}(\sigma,w^{\prime}(v))\leq\phi_{s+t-1,w}(\sigma,w(v))\leq\phi_{w}(w(v))$. For each $j\in[t-1]$, let $d_{j}\coloneqq d_{G}(v,v_{s+j-1})$ and $w_{j}\coloneqq w(v_{s+j-1})$. Again, by Lemma B.1, we have
$f_{(v_{s},v_{s+1},\ldots,v_{s+t-2});w}(w(v))$ $=\left(\prod_{i=1}^{t-1}\frac{d_{i}}{d_{i}+1}\right)w(v)+\sum_{i=1}^{t-1}\left(\prod_{j=i+1}^{t-1}\frac{d_{j}}{d_{j}+1}\right)\frac{1}{d_{i}+1}w_{i}$
$=\left(\prod\limits_{i=1}^{t-1}\frac{d_{i}}{d_{i}+1}\right)w(v)+\frac{\sum\limits_{i=1}^{t-1}\left(\prod\limits_{j=1}^{i-1}(1+d_{j})\right)\left(\prod\limits_{k=i+1}^{t-1}d_{k}\right)w_{i}}{\prod\limits_{i=1}^{t-1}(d_{i}+1)}$
$\geq\left(\prod_{i=1}^{t-1}\frac{d_{i}}{d_{i}+1}\right)w(v)+\frac{\sum\limits_{i=1}^{t-1}\left(\prod\limits_{j=1}^{i-1}(1+d_{j})\right)\left(\prod\limits_{k=i+1}^{t-1}d_{k}\right)}{\prod\limits_{i=1}^{t-1}(d_{i}+1)}\frac{\sum\limits_{i=1}^{t-1}w_{i}}{t-1}$
$=\left(\prod\limits_{i=1}^{t-1}\frac{d_{i}}{d_{i}+1}\right)w(v)+\left(1-\prod\limits_{i=1}^{t-1}\frac{d_{i}}{d_{i}+1}\right)\frac{\sum\limits_{i=1}^{t-1}w_{i}}{t-1}.$
Since $\{d_{1},d_{2},\ldots,d_{t-1}\}$ is an integer interval containing $1$, we have $\prod_{i=1}^{t-1}\frac{d_{i}}{d_{i}+1}\leq\frac{1}{t}$. As $w(v)<w^{\prime}(v)=\frac{1}{t}(w(v)+\sum_{i=1}^{t-1}w_{i})$, we have $w(v)<\frac{\sum_{i=1}^{t-1}w_{i}}{t-1}$. Therefore,
$\left(\prod_{i=1}^{t-1}\frac{d_{i}}{d_{i}+1}\right)w(v)+\left(1-\prod_{i=1}^{t-1}\frac{d_{i}}{d_{i}+1}\right)\frac{\sum\limits_{i=1}^{t-1}w_{i}}{t-1}\geq\frac{1}{t}w(v)+\left(1-\frac{1}{t}\right)\frac{\sum\limits_{i=1}^{t-1}w_{i}}{t-1}=w^{\prime}(v).$
This implies that $f_{(v_{s},v_{s+1},\ldots,v_{s+t-2});w}(w(v))\geq w^{\prime}(v)=f_{(v_{s},v_{s+1},\ldots,v_{s+t-2});w^{\prime}}(w^{\prime}(v))$ and therefore $\phi_{w}(w(v))\geq\phi_{s,w}(\sigma,w(v))\geq\phi_{s,w^{\prime}}(\sigma,w^{\prime}(v))=\phi_{w^{\prime}}(w^{\prime}(v))$, where the last equality follows from Lemma 3.3. ∎
Combining Lemma B.2 with Lemma 3.1Lemma 3.2, and Lemma 3.3 gives the desired proof of Proposition 3.5.